OpenStudy (anonymous):

How do I simplify (X^6)^(2/3)+ (X^6)^(1/3)

OpenStudy (anonymous):

multiple the exponents together ex \[(x^3)^2=x^{(3*2)}\]

OpenStudy (anonymous):

i get x^6 as my answer but when i check it isn't right

OpenStudy (anonymous):

cause it isnt right

OpenStudy (anonymous):

6*2/3=? 6*1/3=?

Directrix (directrix):

@soccernut88 --> To get this answer, i get x^6 as my answer but when i check it isn't right ... did you add x^4 + x^2 to get x^6. If so, that is not correct. It is true that x^4 * x^2 = x^6 but that does not hold for addition of exponential expressions.